The perimeter of a triangle is 136 cm. If the three sides of the triangle are (x+7) cm, (2x-5) cm, and (3x+2) cm, what is the length of each side?
Answer by mananth(16949)

x+7) cm, (2x-5) cm, and (3x+2)
..
(x+7)+(2x-5) + (3x+2)= 136
6x+4 = 136
-4 -4
6x=132
/6
x=22
..
sides
------
(x+7)= 22+7 = 29
(2x-5)=44-5 = 39
(3x+2)=66+2=68
...
m.ananth@hotmail.ca